# Signal attenuation enables scalable decentralized multi-agent reinforcement learning over networks

### v1

@@ III-B Signal Attenuation and Decentralization
      $\leq\left\lVert\lambda\right\rVert_{1}\left[\frac{ML^{i}|\mathcal{N}_{\kappa}^{-1}(i)|}{g^{2}(\kappa,R)}+\varepsilon_{\kappa}\right]$, for all $\lambda\in\mathbb{R}^{n}$.
Proof sketch. As in Theorem 1, the case where $f^{i}=c^{i}$ is trivial due to Assumption 2, so we consider the case where $f^{i}=r^{i}$. Part (i) holds by Theorem 1 and an easy bounding argument leveraging the fact that the weighting function $w_{i}$ in (15) is a convex combination. Part (ii) follows from Theorem 1 by an application of the Cauchy-Schwarz inequality and Assumption 5. Part (iii) follows by considering (ii) over all $\mathcal{N}$, and (iv) follows from (iii) combined with Assumption 6 and an application of Cauchy-Schwarz. $\square$
@@
Theorem 2 provides approximate policy gradient expressions that can be computed using only local neighborhood information and establishes corresponding error bounds depending on the system model, communication radius, choice of $\kappa$, policy design, and placement of radars within the network. The bounds arise due to signal decay inherent in the radar range equations (2)-(3), manifested in the reward (5). These bounds provide a guide to the selection $\kappa$, design of $g$ from Assumption 1, radar placement based on the effective communication range $R$, properties of the system model captured by $M$, properties of the parametric policy class through the Lipschitz constants $L_{i}$, and the number of radars $n$. Importantly, the result demonstrates that, for appropriate choice of design parameters given the underlying system, the expressions presented in Theorem 2 provide good approximations of the desired policy gradients while using only local neighborhood information. Armed with these policy gradient expressions, we next turn to development of decentralized MARL algorithms leveraging them to solve the problems proposed in Section II-B.
@@ IV Algorithms
1:  Input: stepsizes $\alpha_{t},\beta_{t},\eta_{t}$
@@
7:     $\widehat{\mu^{c^{i}}_{t}}=(1-\eta_{t})\widehat{\mu^{c^{i}}_{t-1}}+\eta_{t}c^{i}_{t}$
8:     $\widehat{\mu^{r^{i}}_{t}}=(1-\eta_{t})\widehat{\mu^{r^{i}}_{t-1}}+\eta_{t}r^{i}_{t}$
@@
10:     $\widetilde{Q}^{c^{i}}_{t}=\text{{UPDATE\_Q}}(\widetilde{Q}^{c^{i}}_{t-1},c^{i}_{t},\widehat{\mu^{c^{i}}_{t}},s^{\mathcal{N}_{\kappa}(i)}_{t},a_{t}^{\mathcal{N}_{\kappa}(i)},\eta_{t})$
11:     $\widetilde{Q}^{r^{i}}_{t}=\text{{UPDATE\_Q}}(\widetilde{Q}^{r^{i}}_{t-1},r^{i}_{t},\widehat{\mu^{r^{i}}_{t}},s^{\mathcal{N}_{\kappa}(i)}_{t},a_{t}^{\mathcal{N}_{\kappa}(i)},\eta_{t})$
@@
1:  Input: stepsizes $\alpha_{t},\beta_{t},\eta_{t}$
2:  Initialize: initialize $s_{0},\theta_{0},\lambda_{0},\nu_{0}$, set $t=\widehat{\mu^{c^{i}}_{0}}=\widehat{\mu^{r^{i}}_{0}}=0$, and set all entries of $\widetilde{Q}^{r^{i}},\widetilde{Q}^{c^{i}}$ to 0, for all $i\in\mathcal{N}$
@@
7:     $\widehat{\mu^{c^{i}}_{t}}=(1-\eta_{t})\widehat{\mu^{c^{i}}_{t-1}}+\eta_{t}c^{i}_{t}$
8:     $\widehat{\mu^{r^{i}}_{t}}=(1-\eta_{t})\widehat{\mu^{r^{i}}_{t-1}}+\eta_{t}r^{i}_{t}$
@@
10:     $\widetilde{Q}^{c^{i}}_{t}=\text{{UPDATE\_Q}}(\widetilde{Q}^{c^{i}}_{t-1},c^{i}_{t},\widehat{\mu^{c^{i}}_{t}},s^{\mathcal{N}_{\kappa}(i)}_{t},a_{t}^{\mathcal{N}_{\kappa}(i)},\eta_{t})$
11:     $\widetilde{Q}^{r^{i}}_{t}=\text{{UPDATE\_Q}}(\widetilde{Q}^{r^{i}}_{t-1},r^{i}_{t},\widehat{\mu^{r^{i}}_{t}},s^{\mathcal{N}_{\kappa}(i)}_{t},a_{t}^{\mathcal{N}_{\kappa}(i)},\eta_{t})$
@@
$\theta^{i}_{t+1}$ $=\theta^{i}_{t}-\alpha_{t}\Big{(}\Big{(}1+\sum_{j\in\mathcal{N}_{\kappa}(i)}\nu^{j}\Big{)}\widehat{h^{i}_{c^{i},t}}-\sum_{j\in\mathcal{N}_{\kappa}(i)}\lambda^{j}\widehat{h^{i}_{r^{j},t}}\Big{)}$
$\lambda^{i}_{t+1}$ $=\lambda^{i}_{t}+\beta_{t}\Big{(}\gamma_{min}-\widehat{\mu^{r^{i}}_{t}}\Big{)}$
@@
1:  Input: $\widetilde{Q}^{f^{i}}_{t-1},f^{i}_{t},\widehat{\mu^{f^{i}}_{t}},s_{t}^{\mathcal{N}_{\kappa}(i)},a_{t}^{\mathcal{N}_{\kappa}(i)},\eta_{t}$
@@
$\widetilde{Q}^{f^{i}}_{t}(s_{t}^{\mathcal{N}_{\kappa}(i)},a_{t}^{\mathcal{N}_{\kappa}(i)})$ $=(1-\eta_{t})\widetilde{Q}^{f^{i}}_{t-1}(s_{t}^{\mathcal{N}_{\kappa}(i)},a_{t}^{\mathcal{N}_{\kappa}(i)})$
$+\eta_{t}\left(f^{i}_{t}-\widehat{\mu^{f^{i}}_{t}}+\widetilde{Q}^{f^{i}}_{t-1}(s_{t}^{\mathcal{N}_{\kappa}(i)},a_{t}^{\mathcal{N}_{\kappa}(i)})\right)$
@@ IV-A Sum-of-SINRs Maximization
Our goal is to perform the necessary updates in a decentralized manner, where each agent only needs information available within its local neighborhood. Fortunately, given the approximations provided by Theorem 2, parts (i)-(iii), we have
@@
where $\widehat{\mu^{c^{j}}}(\theta)\approx J_{c^{j}}(\theta)$ is a suitable approximation, such as an empirical average. Using these expressions, we provide the D-SP-PG scheme for solving (II-B1) in Algorithm 1.
@@ IV-B Power Minimization with SINR Threshold
$\mathscr{L}_{R}(\theta,\lambda,\nu)$ $=J_{c}(\theta)+\sum_{j\in\mathcal{N}}\lambda^{j}\left(\gamma_{min}-J_{r^{j}}(\theta)\right)$ (22)
@@
where we recall that $J_{c}(\theta)=\sum_{j\in\mathcal{N}}J_{c^{j}}(\theta)$. In order to solve (II-B2), our goal is to instead solve the saddle point problem $\min_{\theta}\max_{\lambda,\nu}\mathscr{L}_{R}(\theta,\lambda,\nu)$ by alternating between stochastic gradient descent in $\theta$ and ascent in $\lambda$ and $\nu$. Differentiating (22), (23) with respect to $\theta^{i}$ gives
$\nabla_{\theta^{i}}$ $\mathscr{L}_{R}(\theta,\lambda,\nu)=\nabla_{\theta^{i}}J_{c}(\theta)-\sum_{j\in\mathcal{N}}\lambda^{j}\nabla_{\theta^{i}}J_{r^{j}}(\theta)$ (24)
@@
$=\nabla_{\theta^{i}}J_{c^{i}}(\theta)-\sum_{j\in\mathcal{N}}\lambda^{j}\nabla_{\theta^{i}}J_{r^{j}}(\theta)+\sum_{k\in\mathcal{N}_{\kappa}(i)}\nu^{k}\nabla_{\theta^{i}}J_{c^{i}}(\theta)$ (26)
$=\Big{(}1+\sum_{k\in\mathcal{N}_{\kappa}(i)}\nu^{j}\Big{)}\nabla_{\theta^{i}}J_{c^{i}}(\theta)-\sum_{j\in\mathcal{N}}\lambda^{j}\nabla_{\theta^{i}}J_{r^{j}}(\theta),$ (27)
where the first equality holds by Assumption 2 combined with our assumption that the $\kappa$-hop neighbor relation is symmetric, i.e., that $i\in\mathcal{N}_{\kappa}(j)$ if and only if $j\in\mathcal{N}_{\kappa}(i)$. Finally, differentiating with respect to $\lambda^{i}$ and $\nu^{i}$ yields
$\nabla_{\lambda^{i}}\mathscr{L}_{R}(\theta,\lambda,\nu)$ $=\gamma_{min}-J_{r^{i}}(\theta),$ (28)
$\nabla_{\nu^{i}}\mathscr{L}_{R}(\theta,\lambda,\nu)$ $=u^{i}-\sum_{j\in\mathcal{N}_{\kappa}(i)}J_{c^{j}}(\theta).$ (29)
Our goal is again to perform the necessary updates in a decentralized manner, with each agent only using locally available information. Using the approximations provided by Theorem 2, parts (i)-(iv), we have
$\nabla_{\theta^{i}}\mathscr{L}_{R}(\theta,\lambda,\nu)$ $\approx\Big{(}1+\sum_{k\in\mathcal{N}_{\kappa}(i)}\nu^{k}\Big{)}\widehat{h^{i}_{c^{i}}}(\theta)-\sum_{j\in\mathcal{N}_{\kappa}(i)}\lambda^{j}\widehat{h^{i}_{r^{j}}}(\theta)$ (30)
$\nabla_{\lambda^{i}}\mathscr{L}_{R}(\theta,\lambda,\nu)$ $\approx\gamma_{min}-\widehat{\mu^{r^{i}}}(\theta)$ (31)
$\nabla_{\nu^{i}}\mathscr{L}_{R}(\theta,\lambda,\nu)$ $\approx u^{i}-\sum_{j\in\mathcal{N}_{\kappa}(i)}\widehat{\mu^{c^{j}}}(\theta),$ (32)

### v2

@@ III-B Signal Attenuation and Decentralization
      $\leq\sum_{j\in\mathcal{N}_{\kappa}(i)}\left\lvert\eta^{j}\right\rvert\frac{ML^{i}\left\lvert\mathcal{N}_{\kappa}^{-1}(j)\right\rvert}{g^{2}(\kappa,R)}+\sum_{j\in\mathcal{N}_{\kappa}^{-1}(i)}\left\lvert\eta^{j}\right\rvert\varepsilon_{\kappa}$, for all $\eta\in\mathbb{R}^{n}$.
Proof.
As in Theorem 1, the $f^{i}=c^{i}$ case is trivial by Assumption 2, so consider only the $f^{i}=r^{i}$ case. Fix $w^{i},s\in\mathcal{S}$, and $a\in\mathcal{A}$.
Part (i). By the definition of $\widetilde{Q}^{r^{i}}_{\theta}$, we have
$\left\lvert\widetilde{Q}^{r^{i}}_{\theta}(s^{\mathcal{N}_{\kappa}(i)},a^{\mathcal{N}_{\kappa}(i)})-Q^{r^{i}}_{\theta}(s,a)\right\rvert$ (52)
$\leq\sum_{\bar{s}\in\mathcal{S}^{\mathcal{N}_{\kappa}^{-1}(i)},\bar{a}\in\mathcal{A}^{\mathcal{N}_{\kappa}^{-1}(i)}}w^{i}(\bar{s},\bar{a})\left\lvert Q^{r^{i}}_{\theta}\left((s^{\mathcal{N}_{\kappa}(i)},\bar{s}),(a^{\mathcal{N}_{\kappa}(i)},\bar{a})\right)-Q^{r^{i}}_{\theta}(s,a)\right\rvert$ (53)
$\leq\sum_{\bar{s}\in\mathcal{S}^{\mathcal{N}_{\kappa}^{-1}(i)},\bar{a}\in\mathcal{A}^{\mathcal{N}_{\kappa}^{-1}(i)}}w^{i}(\bar{s},\bar{a})\frac{M\left\lvert\mathcal{N}_{\kappa}^{-1}(i)\right\rvert}{g^{2}(\kappa,R)}=\frac{M\left\lvert\mathcal{N}_{\kappa}^{-1}(i)\right\rvert}{g^{2}(\kappa,R)}.$ (54)
Part (ii). By the policy gradient theorem [18],
$\nabla_{\theta^{i}}J_{r^{j}}(\theta)=\mathbb{E}_{\pi_{\theta}}\left[Q^{r^{j}}_{\theta}(s,a)\nabla_{\theta^{i}}\log\pi^{i}_{\theta^{i}}(a^{i}|s^{i})\right].$ (55)
This means we can write
$\left\lVert\widehat{h^{i}_{r^{j}}}(\theta)-\nabla_{\theta^{i}}J_{r^{j}}(\theta)\right\rVert$ (56)
$=\mathbb{E}_{\pi_{\theta}}\big{[}\big{(}\widetilde{Q}^{r^{j}}_{\theta}(s^{\mathcal{N}_{\kappa}(j)},a^{\mathcal{N}_{\kappa}(j)})-Q^{r^{j}}_{\theta}(s,a)\big{)}\nabla_{\theta^{i}}\log\pi^{i}_{\theta^{i}}(a^{i}|s^{i})\big{]}$ (57)
$\stackrel{{\scriptstyle\textnormal{(h)}}}{{\mathstrut{\leq}}}\mathbb{E}_{\pi_{\theta}}\left[\left\lvert\widetilde{Q}^{r^{j}}_{\theta}(s^{\mathcal{N}_{\kappa}(j)},a^{\mathcal{N}_{\kappa}(j)})-Q^{r^{j}}_{\theta}(s,a)\right\rvert\cdot\left\lVert\nabla_{\theta^{i}}\log\pi^{i}_{\theta^{i}}(a^{i}|s^{i})\right\rVert\right]$ (58)
$\stackrel{{\scriptstyle\textnormal{(i)}}}{{\mathstrut{\leq}}}\mathbb{E}_{\pi_{\theta}}\left[\frac{M\left\lvert\mathcal{N}_{\kappa}^{-1}(j)\right\rvert}{g^{2}(\kappa,R)}\cdot L^{i}\right]=\frac{ML^{i}\left\lvert\mathcal{N}_{\kappa}^{-1}(j)\right\rvert}{g^{2}(\kappa,R)},$ (59)
where (a) follows by Jensen’s inequality and the Cauchy-Schwarz inequality, and (b) follows by Theorem 1 and Assumption 5.
Part (iii). By (15) and the definition of $\widehat{h^{i}_{r}}(\theta)$, we have
$\nabla_{\theta^{i}}J_{r}(\theta)$ $=\sum_{j\in\mathcal{N}}\nabla_{\theta^{i}}J_{r^{j}}(\theta),\quad\widehat{h^{i}_{r}}(\theta)=\sum_{j\in\mathcal{N}_{\kappa}(i)}\widehat{h^{i}_{r^{j}}}(\theta).$ (60)
We thus have
$\left\lVert\widehat{h^{i}_{r}}(\theta)-\nabla_{\theta^{i}}J_{r}(\theta)\right\rVert$ (61)
$=\left\lVert\sum_{j\in\mathcal{N}_{\kappa}(i)}\left(\widehat{h^{i}_{r^{j}}}(\theta)-\nabla_{\theta^{i}}J_{r^{j}}(\theta)\right)-\sum_{j\in\mathcal{N}_{\kappa}^{-1}(i)}\nabla_{\theta^{i}}J_{r^{j}}(\theta)\right\rVert$ (62)
$\leq\sum_{j\in\mathcal{N}_{\kappa}(i)}\left\lVert\widehat{h^{i}_{r^{j}}}(\theta)-\nabla_{\theta^{i}}J_{r^{j}}(\theta)\right\rVert+\sum_{j\in\mathcal{N}_{\kappa}^{-1}(i)}\left\lVert\nabla_{\theta^{i}}J_{r^{j}}(\theta)\right\rVert$ (63)
$\stackrel{{\scriptstyle\textnormal{(j)}}}{{\mathstrut{\leq}}}\sum_{j\in\mathcal{N}_{\kappa}(i)}\frac{ML^{i}\left\lvert\mathcal{N}_{\kappa}^{-1}(j)\right\rvert}{g^{2}(\kappa,R)}+\sum_{j\in\mathcal{N}_{\kappa}^{-1}}\varepsilon_{\kappa}$ (64)
$\stackrel{{\scriptstyle\textnormal{(k)}}}{{\mathstrut{\leq}}}\frac{ML^{i}\bar{n}\left\lvert\mathcal{N}_{\kappa}^{-1}(j)\right\rvert}{g^{2}(\kappa,R)}+\left\lvert\mathcal{N}_{\kappa}^{-1}(i)\right\rvert\varepsilon_{\kappa},$ (65)
where (a) follows by Part (ii) and Assumption 6, and (b) follows by the definition of $\bar{n}$ in the statement of the theorem.
Part (iv). Fix $\eta\in\mathbb{R}^{n}$. We know that
$\left\lVert\sum_{j\in\mathcal{N}_{\kappa}(i)}\eta^{j}\widehat{h^{i}_{r^{j}}}(\theta)-\sum_{j\in\mathcal{N}}\eta^{j}\nabla_{\theta^{i}}J_{r^{j}}(\theta)\right\rVert$ (66)
$=\left\lVert\sum_{j\in\mathcal{N}_{\kappa}(i)}\eta^{j}\left[\widehat{h^{i}_{r^{j}}}(\theta)-\nabla_{\theta^{i}}J_{r^{j}}(\theta)\right]-\sum_{j\in\mathcal{N}_{\kappa}^{-1}(i)}\eta^{j}\nabla_{\theta^{i}}J_{r^{j}}(\theta)\right\rVert$ (67)
$\stackrel{{\scriptstyle\textnormal{(l)}}}{{\mathstrut{\leq}}}\sum_{j\in\mathcal{N}_{\kappa}(i)}\left\lvert\eta^{j}\right\rvert\left\lVert\widehat{h^{i}_{r^{j}}}(\theta)-\nabla_{\theta^{i}}J_{r^{j}}(\theta)\right\rVert+\sum_{j\in\mathcal{N}_{\kappa}^{-1}(i)}\left\lvert\eta^{j}\right\rvert\left\lVert\nabla_{\theta^{i}}J_{r^{j}}(\theta)\right\rVert$ (68)
$\stackrel{{\scriptstyle\textnormal{(m)}}}{{\mathstrut{\leq}}}\sum_{j\in\mathcal{N}_{\kappa}(i)}\left\lvert\eta^{j}\right\rvert\frac{ML^{i}\left\lvert\mathcal{N}_{\kappa}^{-1}(j)\right\rvert}{g^{2}(\kappa,R)}+\sum_{j\in\mathcal{N}_{\kappa}^{-1}(i)}\left\lvert\eta^{j}\right\rvert\varepsilon_{\kappa},$ (69)
where (a) follows by the triangle and Cauchy-Schwarz inequalities and (b) follows by Part (ii) and Assumption 6. ∎
@@
Theorem 2 provides approximate policy gradient expressions that can be computed using only local neighborhood information and establishes corresponding error bounds depending on the system model, communication radius, choice of $\kappa$, policy design, and placement of radars within the network. The bounds arise due to signal decay inherent in the radar range equations (2)-(3), manifested in the reward (5). These bounds provide a guide to the selection of $\kappa$, design of $g$ from Assumption 1, radar placement based on the effective communication range $R$, properties of the system model captured by $M$, properties of the parametric policy class through the Lipschitz constants $L_{i}$, and the number of radars $n$. Importantly, the result demonstrates that, for appropriate choice of design parameters given the underlying system, the expressions presented in Theorem 2 provide good approximations of the desired policy gradients while using only local neighborhood information. Armed with these policy gradient expressions, we next turn to development of decentralized MARL algorithms leveraging them to solve the problems proposed in Section II-B.
@@ IV Algorithms
1:  Input: stepsizes $\alpha_{t},\beta_{t},\zeta_{t}$
@@
7:     $\widehat{\mu^{c^{i}}_{t}}=(1-\zeta_{t})\widehat{\mu^{c^{i}}_{t-1}}+\zeta_{t}c^{i}_{t}$
8:     $\widehat{\mu^{r^{i}}_{t}}=(1-\zeta_{t})\widehat{\mu^{r^{i}}_{t-1}}+\zeta_{t}r^{i}_{t}$
@@
10:     $\widetilde{Q}^{c^{i}}_{t}=\text{{UPDATE\_Q}}(\widetilde{Q}^{c^{i}}_{t-1},c^{i}_{t},\widehat{\mu^{c^{i}}_{t}},s^{\mathcal{N}_{\kappa}(i)}_{t},a_{t}^{\mathcal{N}_{\kappa}(i)},\zeta_{t})$
11:     $\widetilde{Q}^{r^{i}}_{t}=\text{{UPDATE\_Q}}(\widetilde{Q}^{r^{i}}_{t-1},r^{i}_{t},\widehat{\mu^{r^{i}}_{t}},s^{\mathcal{N}_{\kappa}(i)}_{t},a_{t}^{\mathcal{N}_{\kappa}(i)},\zeta_{t})$
@@
1:  Input: stepsizes $\alpha_{t},\beta_{t},\zeta_{t}$
2:  Initialize: initialize $s_{0},\theta_{0},\eta_{0},\nu_{0}$, set $t=\widehat{\mu^{c^{i}}_{0}}=\widehat{\mu^{r^{i}}_{0}}=0$, and set all entries of $\widetilde{Q}^{r^{i}},\widetilde{Q}^{c^{i}}$ to 0, for all $i\in\mathcal{N}$
@@
7:     $\widehat{\mu^{c^{i}}_{t}}=(1-\zeta_{t})\widehat{\mu^{c^{i}}_{t-1}}+\zeta_{t}c^{i}_{t}$
8:     $\widehat{\mu^{r^{i}}_{t}}=(1-\zeta_{t})\widehat{\mu^{r^{i}}_{t-1}}+\zeta_{t}r^{i}_{t}$
@@
10:     $\widetilde{Q}^{c^{i}}_{t}=\text{{UPDATE\_Q}}(\widetilde{Q}^{c^{i}}_{t-1},c^{i}_{t},\widehat{\mu^{c^{i}}_{t}},s^{\mathcal{N}_{\kappa}(i)}_{t},a_{t}^{\mathcal{N}_{\kappa}(i)},\zeta_{t})$
11:     $\widetilde{Q}^{r^{i}}_{t}=\text{{UPDATE\_Q}}(\widetilde{Q}^{r^{i}}_{t-1},r^{i}_{t},\widehat{\mu^{r^{i}}_{t}},s^{\mathcal{N}_{\kappa}(i)}_{t},a_{t}^{\mathcal{N}_{\kappa}(i)},\zeta_{t})$
@@
$\theta^{i}_{t+1}$ $=\theta^{i}_{t}-\alpha_{t}\Big{(}\Big{(}1+\sum_{j\in\mathcal{N}_{\kappa}(i)}\nu^{j}\Big{)}\widehat{h^{i}_{c^{i},t}}-\sum_{j\in\mathcal{N}_{\kappa}(i)}\eta^{j}\widehat{h^{i}_{r^{j},t}}\Big{)}$
$\eta^{i}_{t+1}$ $=\eta^{i}_{t}+\beta_{t}\Big{(}\gamma_{min}-\widehat{\mu^{r^{i}}_{t}}\Big{)}$
@@
1:  Input: $\widetilde{Q}^{f^{i}}_{t-1},f^{i}_{t},\widehat{\mu^{f^{i}}_{t}},s_{t}^{\mathcal{N}_{\kappa}(i)},a_{t}^{\mathcal{N}_{\kappa}(i)},\zeta_{t}$
@@
$\widetilde{Q}^{f^{i}}_{t}(s_{t}^{\mathcal{N}_{\kappa}(i)},a_{t}^{\mathcal{N}_{\kappa}(i)})$ $=(1-\zeta_{t})\widetilde{Q}^{f^{i}}_{t-1}(s_{t}^{\mathcal{N}_{\kappa}(i)},a_{t}^{\mathcal{N}_{\kappa}(i)})$
$+\zeta_{t}\left(f^{i}_{t}-\widehat{\mu^{f^{i}}_{t}}+\widetilde{Q}^{f^{i}}_{t-1}(s_{t}^{\mathcal{N}_{\kappa}(i)},a_{t}^{\mathcal{N}_{\kappa}(i)})\right)$
@@ IV-A Sum-of-SINRs Maximization
We can now present a decentralized learning algorithm for approximately solving this problem, where each agent only needs information available within its local $\kappa$-hop neighborhood. Using the approximations provided by Theorem 2, for suitable choice of $\kappa$ and $R$ we have
@@
where $\widehat{\mu^{c^{j}}}(\theta)\approx J_{c^{j}}(\theta)$ is a suitable approximation, such as a cumulative or exponential moving average. Using these expressions, we provide the D-SP-PG scheme for solving (II-B1) in Algorithm 1.
@@ IV-B Power Minimization with SINR Threshold
$\mathscr{L}_{R}(\theta,\eta,\nu)$ $=J_{c}(\theta)+\sum_{j\in\mathcal{N}}\eta^{j}\left(\gamma_{min}-J_{r^{j}}(\theta)\right)$ (76)
@@
where we recall that $J_{c}(\theta)=\sum_{j\in\mathcal{N}}J_{c^{j}}(\theta)$. In order to solve (II-B2), we instead solve the saddle point problem $\min_{\theta}\max_{\eta,\nu}\mathscr{L}_{R}(\theta,\eta,\nu)$ by alternating between stochastic gradient descent in $\theta$ and ascent in $\eta$ and $\nu$. Differentiating (76), (77) with respect to $\theta^{i}$ gives
$\nabla_{\theta^{i}}$ $\mathscr{L}_{R}(\theta,\eta,\nu)=\nabla_{\theta^{i}}J_{c}(\theta)-\sum_{j\in\mathcal{N}}\eta^{j}\nabla_{\theta^{i}}J_{r^{j}}(\theta)$ (78)
@@
$=\nabla_{\theta^{i}}J_{c^{i}}(\theta)-\sum_{j\in\mathcal{N}}\eta^{j}\nabla_{\theta^{i}}J_{r^{j}}(\theta)+\sum_{k\in\mathcal{N}_{\kappa}(i)}\nu^{k}\nabla_{\theta^{i}}J_{c^{i}}(\theta)$ (80)
$=\Big{(}1+\sum_{k\in\mathcal{N}_{\kappa}(i)}\nu^{j}\Big{)}\nabla_{\theta^{i}}J_{c^{i}}(\theta)-\sum_{j\in\mathcal{N}}\eta^{j}\nabla_{\theta^{i}}J_{r^{j}}(\theta),$ (81)
where the first equality holds by Assumption 2 combined with our assumption that the $\kappa$-hop neighbor relation is symmetric, i.e., that $i\in\mathcal{N}_{\kappa}(j)$ if and only if $j\in\mathcal{N}_{\kappa}(i)$. Finally, differentiating with respect to $\eta^{i}$ and $\nu^{i}$ yields
$\nabla_{\eta^{i}}\mathscr{L}_{R}(\theta,\eta,\nu)$ $=\gamma_{min}-J_{r^{i}}(\theta),$ (82)
$\nabla_{\nu^{i}}\mathscr{L}_{R}(\theta,\eta,\nu)$ $=u^{i}-\sum_{j\in\mathcal{N}_{\kappa}(i)}J_{c^{j}}(\theta).$ (83)
Our goal is again to perform the necessary updates in a decentralized manner, with each agent only using locally available information. Using the approximations provided by Theorem 2, for suitable choice of $\kappa$ and $R$ we have
$\nabla_{\theta^{i}}\mathscr{L}_{R}(\theta,\eta,\nu)$ $\approx\Big{(}1+\sum_{k\in\mathcal{N}_{\kappa}(i)}\nu^{k}\Big{)}\widehat{h^{i}_{c^{i}}}(\theta)-\sum_{j\in\mathcal{N}_{\kappa}(i)}\eta^{j}\widehat{h^{i}_{r^{j}}}(\theta)$ (84)
$\nabla_{\eta^{i}}\mathscr{L}_{R}(\theta,\eta,\nu)$ $\approx\gamma_{min}-\widehat{\mu^{r^{i}}}(\theta)$ (85)
$\nabla_{\nu^{i}}\mathscr{L}_{R}(\theta,\eta,\nu)$ $\approx u^{i}-\sum_{j\in\mathcal{N}_{\kappa}(i)}\widehat{\mu^{c^{j}}}(\theta),$ (86)
@@
Remark 2.
We note that, in order to lay the groundwork for convergence analysis of our algorithms in future work, we have presented the $Q$ updates of Algorithm 3 in the easily analyzable tabular form, which is only applicable to finite state and action spaces. We emphasize that practical variants leveraging standard neural network architectures for $Q$ function approximation can be substituted in Algorithm 3 to enable application to the continuous spaces of Section II.